# TwinDiffusion: Enhancing Coherence and Efficiency in Panoramic Image Generation with Diffusion Models

### v1

@@ 1 Introduction
Over the past few years, diffusion models [41, 16, 28] have demonstrated their creativity in generation tasks. They define a pair of forward and reverse Markov chains to learn the data distribution, which bypasses the limitations in other types of generative models like GANs [7, 4], VAEs [23, 6] and Flows [19]. Diffusion models have risen as effective tools for broad applications, spanning from high-quality images to multi-type content creation [5, 29, 38, 14]. With the growing interest in full information records, immersive virtual reality, artistic expression, and historical preservation, the synthesis of long scrolls is gaining more and more attention, particularly in panoramic image generation tasks [2, 11, 42].
@@ 2 Related Work
Diffusion models [34] are inspired by non-equilibrium thermodynamics. They define two Markov chains for forward and reverse processes, namely diffusion and denoising. The forward process is to perturb a data distribution $x_{0}\sim q(x)$ into a standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$ with $T$ steps of noise injection. This process is reversed to recreate the sample $x_{0}$ that obeys the original data distribution from a Gaussian noise input. Specifically, the reverse process involves training a network to approximate $q(x_{t-1}\mid x_{t})$, and then sampling from $\mathcal{N}(0,I)$ iteratively with the trained model. From DDPM [16] to DDIM [35] to LDM [28], diffusion models have paved the way for text-to-image generation, boosting AI-painting applications like Stable Diffusion [28] and DALLE2 [27]. By capturing the spatial-temporal distribution features, they also show promising prospects in generating multi-modal contents such as videos [3, 10, 21], audio [17, 20], and 3D objects [25, 36].
@@ 3.1 Preliminary
To start, we introduce a pre-trained diffusion model denoted by $\Phi$, operating in a latent space $Z=\mathbb{R}^{h\times w\times c}$ and a textural condition $C$. Employing the deterministic DDIM sampling:
@@ 3.3 Cross Sampling
where $c$ controls the interleaving frequency and $k$ denotes the sampling mode. Staggering in $c$ times, our method dynamically formes a set $\mathcal{X}=\{\mathbf{x}_{i,j}\mid 1\leq i\leq c,1\leq j\leq n\}$ consists of $c$ groups, each containing $n$ overlapping crops, with the spatial locations incrementing by $s_{c}$. Then, in $T$ rounds of denoising, we alternate between using these crop groups for sampling in $c$ staggered spaces.
@@ 4.1 Panorama with Twin Crops
Beyond a single pair of images, we generalize this approach to a sequence of images, i.e., crops $\mathbf{x}_{t}^{1},\mathbf{x}_{t}^{2},\ldots,\mathbf{x}_{t}^{n}$ within panoramas. Our optimization is applied to each pair of neighboring crops, promoting a high degree of similarity between $[\mathbf{x}_{t}^{i}]_{r}$ and $[\mathbf{x}_{t}^{i+1}]_{l}$, thereby resulting in higher-quality panoramas with consecutive twin crops. Refer to Appendix A for more implementation examples.
@@ 4.3 Ablation
In the initial attempts at twin-image generation, we apply the Crop Fusion method throughout the entire diffusion process. However, the outcomes are unsatisfactory as the generated images exhibit a distinct left-right fragmentation, where an appropriate optimization time window is needed to achieve the desirable results. We know that an earlier guidance plays a greater role in the diffusion trajectory. Thus, we decide to confine the optimization period to the early stages of the sampling process. That is, the Crop Fusion is carried out from $t:T\rightarrow{\tau}$ and then stopped during $t:\tau\rightarrow 0$, where $T$ represents the total timestep according to the specific scheduler. The relationship between $\tau$ and the stitching effect of twin images is depicted in Fig. 6. As $\tau$ increases, we can see a gradual transition from under-optimization to over-optimization, with the best results attained by $\tau=T/2$.
Figure 6: The analysis of the proper timestep $\tau$ for introducing the Crop Fusion stage. There’s a progressive stitching effect on $I_{2}^{*}$ with the increase of optimization timestep $\tau$. (i) When $\tau<T/2$, it is under-optimized, as $[I_{2}^{*}]_{l}$ differs significantly from $[I_{1}]_{r}$. (ii) When $\tau=T/2$, it is well-optimized, generating the most natural and seamless $I_{2}^{*}$. (iii) When $\tau>T/2$, it is over-optimized, $I_{2}^{*}$ stays too close with $[I_{1}]_{r}$ while failing to fuse with $[I_{2}^{*}]_{r}$. The bottom lines offer two additional examples that further support our findings.
@@
Figure 7: Further explorations about adjacent control factor $\lambda$. It presents a comparative analysis across a wide range of $\lambda$ values, ranging from 0.1 to 100, to examine their impact on the alignment behavior and visual coherence in panoramic images based on the LPIPS-DISTS weighted metric. The results demonstrate that our approach is robust, with the optimal $\lambda$ value around 1 achieving the best balance between reconciling with adjacent crops and maintaining self-alignment. This is reasonable and in keeping well with our objective.
@@
In particular, we demonstrate the superiority of our method from three aspects. (i) For quality: under the same view stride, our method, with single Crop Fusion or Cross Sampling, both reach a lower LPIPS-DISTS weighted score than the MultiDiffusion, and full TwinDiffusion reaches even lower scores, with the effects of both methods add up. (ii) For efficiency: analyzing the blue and green lines, we can observe that our Cross Sampling approach successfully enables generating comparable results within a fraction of the time required by the original MultiDiffusion, specifically achieving a reduction of N-fold when using a $N\times s_{v}$ value. (iii) For quality-efficiency trade-off: our method takes greater advantage of the efficiency gained from larger $s_{v}$ while maintaining high image quality, striking an optimal balance between quality and speed for panoramic image generation.
@@
Table 1: Here are the findings from our further exploration of the Cross Sampling strategy. Different $s_{c}$ values were tested on all sides, with the best and the second-best results marked in bold and underlined respectively. The better effects are more likely to be achieved when $s_{c}=s_{v}/2,\,s_{v}/3$, meaning that excessively fine cross strides are unnecessary. In the extreme case of $s_{c}=s_{v}/s_{v}=1$, the effect regresses to no Cross Sampling or even worse.
@@ 5 Conclusion
Although TwinDiffusion works well in most cases, it still faces some limitations. Our approach primarily focuses on optimizing the local similarity of adjacent and overlapping image areas, forming a series of image crops with successive content. However, it could not ensure stability in perceiving the overall layout of the images, which may lead to the generation of visually coherent but spatially illogical panoramas. As for potential negative impact, image generation models may involve personal copyright or generate fake, offensive, discriminatory results. Further research should prioritize the responsible use of the relevant technology to avoid generating content in any harmful way.
@@
Appendix A More Implementation Examples in Twin Image Gerenation
Our experimental results of twin images are provided in Fig. A1.
@@ Appendix A More Implementation Examples in Twin Image Gerenation
Appendix B Additional Qualitative Comparison Results with Panoramas
Fig. A2 gives more qualitative comparisons with MultiDiffusion in panoramic image generation.
Figure A2: More qualitative comparison with MultiDiffusion in panorama generation across various text prompts. We mark the areas where improvements have been made.

### v3

@@ 1 Introduction
Over the past few years, diffusion models [39, 16, 27] have demonstrated their creativity in generation tasks. They define a pair of forward and reverse Markov chains to learn the data distribution, which bypasses the limitations in other types of generative models like GANs [7, 4], VAEs [23, 6] and Flows [19]. Diffusion models have risen as effective tools for broad applications, spanning from high-quality images to multi-type content creation [14]. With the growing interest in full information records, immersive virtual reality, artistic expression, and historical preservation, the synthesis of long scrolls is gaining more and more attention, particularly in panoramic image generation tasks [2, 11, 40].
@@ 2 Related Work
Diffusion models [32] are inspired by non-equilibrium thermodynamics. They define two Markov chains for forward and reverse processes, namely diffusion and denoising. The forward process is to perturb a data distribution $x_{0}\sim q(x)$ into a standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$ with $T$ steps of noise injection. This process is reversed to recreate the sample $x_{0}$ that obeys the original data distribution from a Gaussian noise input. Specifically, the reverse process involves training a network to approximate $q(x_{t-1}\mid x_{t})$, and then sampling from $\mathcal{N}(0,I)$ iteratively with the trained model. From DDPM [16] to DDIM [33] to LDM [27], diffusion models have paved the way for text-to-image generation, boosting AI-painting applications like Stable Diffusion [27] and DALLE2 [26]. By capturing the spatial-temporal distribution features, they also show promising prospects in generating multi-modal contents such as videos [3, 10, 21], audio [17, 20], and 3D objects [5, 36, 34].
@@ 3.1 Preliminary
To start with, we introduce a pre-trained diffusion model denoted by $\Phi$, operating in a latent space $Z=\mathbb{R}^{h\times w\times c}$ and a textual condition $C$. Employing the deterministic DDIM sampling:
@@ 3.3 Cross Sampling
where $c$ controls the interleaving frequency and $k$ denotes the sampling mode. Staggering in $c$ times, our method dynamically forms a set $\mathcal{X}=\{\mathbf{x}_{i,j}\mid 1\leq i\leq c,1\leq j\leq n\}$ consists of $c$ groups, each containing $n$ overlapping crops, with spatial locations incrementing by the cross stride $s_{c}$. Then, in $T$ rounds of denoising, we alternate between using these crop groups for sampling in $c$ staggered spaces.
@@ 4.1 Panorama with Twin Crops
Beyond a single pair of images, we generalize this approach to a sequence of images, i.e., crops $\mathbf{x}_{t}^{1},\mathbf{x}_{t}^{2},\ldots,\mathbf{x}_{t}^{n}$ within panoramas. Our optimization is applied to each pair of neighboring crops, promoting a high degree of similarity between $[\mathbf{x}_{t}^{i}]_{r}$ and $[\mathbf{x}_{t}^{i+1}]_{l}$, thereby resulting in higher-quality panoramas with consecutive twin crops as depicted in Fig. 1. Refer to Appendix A for more implementation examples.
@@ 4.3 Ablation
In the initial attempts at twin-image generation, we apply the Crop Fusion method throughout the entire denoising process. However, the outcomes are unsatisfactory as the generated images exhibit a distinct left-right fragmentation, where an appropriate optimization time window is needed to achieve the desirable results. We know that an earlier guidance plays a greater role in the diffusion trajectory. Thus, we decide to confine the optimization period to the early stages of the sampling process. That is, the Crop Fusion is carried out from $t:T\rightarrow{\tau}$ and then stopped during $t:\tau\rightarrow 0$, where $T$ represents the total timestep according to the specific scheduler. The relationship between $\tau$ and the stitching effect of twin images is depicted in Fig. 6. As $\tau$ decreases, we can see a gradual transition from under-optimization to over-optimization, with the best results attained by $\tau=T/2$.
Figure 6: The analysis of the proper timestep $\tau$ for introducing the Crop Fusion stage. There’s a progressive effect on $I_{2}^{*}$ with the decrease of optimization timestep $\tau$. (i) When $\tau>T/2$, it is under-optimized, as $[I_{2}^{*}]_{l}$ differs significantly from $[I_{1}]_{r}$. (ii) When $\tau=T/2$, it is well-optimized, generating the most natural and seamless $I_{2}^{*}$. (iii) When $\tau<T/2$, it is over-optimized, $I_{2}^{*}$ stays too close with $[I_{1}]_{r}$ while failing to fuse with $[I_{2}^{*}]_{r}$. The bottom lines offer two additional examples that further support our findings.
@@
Figure 7: Further explorations about adjacent control factor $\lambda$. It presents a comparative analysis across a wide range of $\lambda$ values, ranging from 0.1 to 100, to examine their impact on the alignment behavior and visual coherence in panoramic images based on the LPIPS-DISTS weighted metric. The results demonstrate that our approach is robust, and with the optimal $\lambda$ value around 1 achieving the best balance between reconciling with adjacent crops and maintaining self-alignment. This is reasonable and in keeping well with our objective.
@@
In particular, it demonstrates its superiority from these three aspects. (i) For quality: under the same view stride, our method, with single Crop Fusion or Cross Sampling, both reach a lower LPIPS-DISTS weighted score than the MultiDiffusion, and full TwinDiffusion reaches even lower scores, with the effects of both methods add up. (ii) For efficiency: analyzing the blue and green lines, we can observe that our Cross Sampling approach successfully enables generating comparable results within a fraction of the time required by the original MultiDiffusion, specifically achieving a reduction of N-fold when using a $N\times s_{v}$ value. (iii) For quality-efficiency trade-off: our method takes greater advantage of the efficiency gained from larger $s_{v}$ while maintaining high image quality, striking an optimal balance between quality and speed for panoramic image generation.
@@
Table 1: Here are the findings from our further exploration of the Cross Sampling strategy. Different $s_{c}$ values were tested on all sides, with the best and the second-best results marked in bold and underlined respectively. The better effects are more likely to be achieved when $s_{c}=s_{v}/2,\,s_{v}/3$, indicating that excessively fine cross strides are unnecessary. In the extreme case of $s_{c}=s_{v}/s_{v}=1$, the effect regresses to no Cross Sampling or even worse.
@@ 5 Conclusion
Although TwinDiffusion works well in most cases, it still faces some limitations. Our approach primarily focuses on optimizing the local similarity of the image areas. However, it could not ensure stability in perceiving the overall layout of the images, which may lead to the generation of visually coherent but spatially illogical panoramas. As for potential negative impact, image generation models may involve personal copyright or generate fake, offensive, discriminatory results. Further research should prioritize the responsible use of the relevant technology to avoid generating content in any harmful way.
@@
Appendix A More Implementation Examples on Twin-Image Generation
Our experimental results of twin images are provided in Fig. A1, meaning the high stability of our method.
@@ Appendix A More Implementation Examples on Twin-Image Generation
Appendix B More Results of Qualitative Comparison on Panorama Generation
Fig. A2 gives more qualitative comparisons with MultiDiffusion on panoramic image generation, highlighting the areas where improvements have been made.
Figure A2: Additional qualitative comparison results on panorama generation with various text prompts.